# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ V.2 Late-time: thermalization via instantons
Due to the explicitly broken time-translation symmetry, the Floquet Hamiltonian is only well defined modulo $\Omega$. However, at $\lambda=0$ the initial Hamiltonian $H_{0}(0)$ can have a many-body bandwidth larger then $\Omega$. The instanton solution in Eq. (22) connects these two limits by describing the band-folding dynamics of $H_{0}(\lambda)$; see e.g. the schematic in Fig. 1. During the instanton event, $(E_{n}-E_{m})$ decreased from $\Omega+\tilde{\Omega}$ to $\Omega-\tilde{\Omega}$. If $|\tilde{\Omega}|>\Omega$, there will appear to be an accidental degeneracy at $\lambda=\tilde{\lambda}$. We expect that this degeneracy will be lifted after including the other matrix elements neglected in Eq.(21). Additionally, we note that when $|\Omega-\tilde{\Omega}|>\Omega$, a single instanton event is not enough to fold $(E_{n},E_{m})$ into an energy interval shorther than $\Omega$. In that case, this can trigger subsequent instanton events driven by the matrix element $\braket{n}{H_{1}(\lambda)}{m}$ (note that $m,n$ is exchanged), where now it satisfies the growth criterion $E_{m}-E_{n}>\Omega$. On the contrary, there will be no subsequent instanton events in $\braket{n}{H_{1}(\lambda)}{m}$ nor $\braket{m}{H_{1}(\lambda)}{n}$ if $|\Omega-\tilde{\Omega}|<\Omega$. Therefore, the energy levels of $H_{0}(\lambda)$ are folded step by step in a series of instanton events. When no further instanton events occur, $H_{0}(\lambda)$ has the property that $|E_{m}-E_{n}|<\Omega$ for any pair of eigenstates $m,n$, and at the same time $H_{1}(\lambda)$ will decay exponentially to zero.
VI Outlook
@@
Going beyond conventional Floquet-Magnus high-frequency expansions and numerical exact diagonalization of microscopic Hamiltonians, here we have employed the Floquet flow formalism to analyze the phenomenology of dynamical freezing from a fresh (and universal) perspective. Most interestingly, we find that the bare driven Hamiltonian flows towards the inevitable thermal fixed point via a series of unstable fixed points even at the dynamically frozen points. Freezing itself is associated with an emergent approximate conservation law, that appears at the first prethermal fixed point along the flow. By combining numerical solutions of the flow equations using exact diagonalization and matrix-product operator-based techniques, as well as analytical approximations that are non-perturbative in the inverse drive frequency ($\Omega$), we have shown that freezing is approximate at any finite $\Omega$. Moreover, the violation of the emergent conservation law, quantified by the commutator $[H_{0}(\lambda),\mathcal{Q}]$, vanishes as $1/\Omega^{2}$ in the high-frequency limit.
@@
One of the key contributions of this work is the analysis of thermalization dynamics of the system beyond the prethermal fixed point, where we find that the flow between the intermediate fixed points can be described using instanton events. For the sake of concreteness, we have chosen to study the driven spin chain as our main numerical example using complementary methods. However, the demonstrated features for freezing and thermalization are quite generic, as demonstrated by the beautiful agreement between our numerical results and the general model-independent analytical computations.
@@
We conclude by discussing several future directions. In the current work, we have only considered the cosine wave drive for simplicity, as there are just two Fourier components in the flow equation. Recently, there has been numerical evidence suggesting that dynamical freezing is more robust for square-wave drive [45, 53]. The flow formalism can be generalized to describe the square-wave drive, by introducing an infinite series of Hamiltonian components corresponding to the Fourier decomposition of the drive function [63]. Whether the flow away from the first prethermal fixed point is similar in character to the example studied here, but with possibly delayed timescales, is an interesting open question.
@@ Acknowledgements.
DC thanks Vedika Khemani for useful discussions, and Dan Mao for related collaborations. We thank Zihao Qi for discussions regarding the isospectral nature of the many-body spectrum during the flow. RM, HG and DC are supported in part by a New Frontier Grant awarded by the College of Arts and Sciences at Cornell University, and by a Sloan research fellowship from the Alfred P. Sloan foundation to DC. HG is also supported by a Bethe-Wilkins-KIC postdoctoral fellowship at Cornell University.
@@ Appendix A Connection between flow-renormalization and Schrieffer-Wolff transformation
Within the flow equation approach [58, 59, 60, 56], the Hamiltonian $H$ is transformed continuously by defining a family of Hamiltonians $H(l)$, where $H(0)=H$ and $H(l)$ flows toward a diagonal (or block-diagonal) form as $l\to\infty$. The flow is governed by the differential equation,
@@
The ideas discussed above can be generalized to the Floquet context using the Sambe representation [87, 63]. In this representation, the time-dependent Hamiltonian $H(t)=\sum_{m}e^{im\Omega t}H_{m}$ is represented in an extended Hilbert space as
@@
Following Ref. [63], the Floquet flow is constructed as a unitary flow in the Sambe representation, which automatically preserves the eigenspectrum of ${\mathcal{H}}$ and hence the spectrum of $H_{F}$. The generator $\eta$ for the flow is
$\eta(l)=\sum_{m}\sum_{M>0}\left[H_{M}(l)\otimes\ket{m+M}\bra{m}-h.c.\right]\,.$ (29)
@@
Appendix B Spectral analysis and quasi-energy eigenvalues
The generators for evolving the Floquet flow equations are constructed to ensure that the evolution retains its isospectral character, preserving the quasienergy provided that both $H_{0}(\lambda)$ and $H_{1}(\lambda)$ are retained [63] . Here, we demonstrate that the flow produces a good approximation for $H_{F}$ in the sense that $H_{0}(\lambda_{c})$ itself is approximately isopectral to $H_{F}$. To verify this numerically, we calculate the quasienergies ( $\epsilon_{n}^{\text{exact}}$) directly from the Floquet unitary [66] using exact diagonalization. We then compute the eigenvalues of $H_{0}(\lambda_{c})$, denoted as $\epsilon_{n}^{H_{0}(\lambda_{c})}$. We define the relative errors,
$\delta=\bigg{\lvert}\frac{\epsilon_{n}^{\text{exact}}-\epsilon_{n}^{H_{0}(\lambda_{c})}}{\epsilon_{n}^{\text{exact}}}\bigg{\rvert}\,$ (30)
whose distribution is plotted in Fig. 8 for increasing system size. We note that for our numerical computations, $\lambda_{c}$ was set equal to $J^{-1}$ with an RK4 step size of $\lambda_{n+1}-\lambda_{n}=0.001$ (in units of $J^{-1}$). We find that this relative error is peaked around $\sim 10^{-5}$. The residual error arises primarily from the inaccuracies introduced by the RK4 step. Furthermore, larger $\Omega$ result in smaller errors; for a fixed $\lambda_{c}$, the norm of $H_{1}(\lambda_{c})$ decreases with increasing $\Omega$, making $H_{0}(\lambda_{c})$ increasingly effective in approximating the actual evolution. This shows that $H_{0}(\lambda_{c})$ is an excellent Hamiltonian to capture the exact many-body dynamics.
Figure 8: Histogram plot of the relative error in quasienergy, $\delta$ (Eq. (30)) as a function of varying system size. The relevant parameters are $J=1$, $J_{2}=0.2$, $\Omega=10$.
Appendix C Early-time analysis
@@
C.1 Review of Floquet-Magnus expansion
@@
$H_{\text{mov}}(t)=W(t)^{\dagger}H(t)W(t)=\exp\left(i\frac{2\sin(\Omega t){\mathrm{Ad}_{H_{1}}}}{\Omega}\right)H_{0}\,.$ (33)
@@
and $I_{00}^{(1)}=-(i\pi)/(2\Omega)$. Since $[h_{m},h_{-m}]=0$, we only need $I_{m0}^{(1)}$ and $I_{0m}^{(1)}$. Therefore we obtain
@@
C.2 Solution of flow equation
@@
[…]
We emphasize that all of the manipulations thus far are exact. To make connection with the Floquet-Magnus expansion above, we can expand Eq. (C.2) in powers of $H_{0}(0)$, i.e.
@@
  • First order term: We first solve for $H_{0}^{(1)}$, which is obtained by setting $\bar{H}_{1}(\lambda)=\bar{H}_{1}(\lambda)^{\dagger}=H_{1}(0)$ in Eq. (C.2), and the associated linear second-order differential equation:
@@
    -- $J_{0}\bigg{(}\frac{A}{\Omega}\bigg{)}Y_{1}\bigg{(}\frac{A}{\Omega}e^{-\Omega\lambda}\bigg{)}\bigg{]}.$
@@
  • Second order term: The term, $H_{0}^{(2)}(\lambda)$, can be obtained by expanding Eq. (C.2) to ${\cal{O}}(\zeta^{2})$, with a resulting equation:
@@
    The source term, $S^{(2)}(\lambda)$, is obtained by expanding the last term of Eq. (C.2) to ${\cal{O}}(\zeta^{2})$,
@@
    +$+\bigg{\{}{\mathrm{Ad}}_{\bar{H}_{1}^{(0)}(\lambda)},{\mathrm{Ad}}_{\bar{H}_{1}^{(1)}(\lambda)^{\dagger}}\bigg{\}}\bigg{]}H_{0}^{(1)}(\lambda)\,.$ (52)
@@
    Interestingly, due to conjugation, $\bar{H}_{1}^{(1)}(\lambda)^{\dagger}=-\bar{H}_{1}^{(1)}(\lambda)$, the two terms in Eq. (C.2) cancel, and hence the source term, $S^{(2)}(\lambda)=0$, in Eq. C.2 vanishes. Combining with the initial condition $H_{0}^{(2)}(0)=\partial_{\lambda}H_{0}^{(2)}(0)=0$, the second order term vanishes identically.
@@
Appendix D Estimate of thermalization timescales

### v2

@@ VI.2 Late-time: thermalization via instantons
Due to the explicitly broken time-translation symmetry, the Floquet Hamiltonian is only well defined modulo $\Omega$. However, at $\lambda=0$ the initial Hamiltonian $H_{0}(0)$ can have a many-body bandwidth larger then $\Omega$. The instanton solution in Eq. (26) connects these two limits by describing the band-folding dynamics of $H_{0}(\lambda)$; see e.g. the schematic in Fig. 1. During the instanton event, $(E_{n}-E_{m})$ decreased from $\Omega+\tilde{\Omega}$ to $\Omega-\tilde{\Omega}$. If $|\tilde{\Omega}|>\Omega$, there will appear to be an accidental degeneracy at $\lambda=\tilde{\lambda}$. We expect that this degeneracy will be lifted after including the other matrix elements neglected in Eq.(25). Additionally, we note that when $|\Omega-\tilde{\Omega}|>\Omega$, a single instanton event is not enough to fold $(E_{n},E_{m})$ into an energy interval shorther than $\Omega$. In that case, this can trigger subsequent instanton events driven by the matrix element $\braket{n|H_{1}(\lambda)|m}$ (note that $m,n$ is exchanged), where now it satisfies the growth criterion $E_{m}-E_{n}>\Omega$. On the contrary, there will be no subsequent instanton events in $\braket{n|H_{1}(\lambda)|m}$ nor $\braket{m|H_{1}(\lambda)|n}$ if $|\Omega-\tilde{\Omega}|<\Omega$. Therefore, the energy levels of $H_{0}(\lambda)$ are folded step by step in a series of instanton events. When no further instanton events occur, $H_{0}(\lambda)$ has the property that $|E_{m}-E_{n}|<\Omega$ for any pair of eigenstates $m,n$, and at the same time $H_{1}(\lambda)$ will decay exponentially to zero.
VII Outlook
@@
Going beyond conventional Floquet-Magnus high-frequency expansions and numerical exact diagonalization of microscopic Hamiltonians, here we have employed the Floquet flow formalism to analyze the phenomenology of dynamical freezing from a fresh (and universal) perspective. Most interestingly, we find that the bare driven Hamiltonian flows through a complex landscape of (unstable) fixed points and shows a strong tendency towards slow thermalization at the dynamically frozen points. Freezing itself is associated with an emergent approximate conservation law, that appears at the first prethermal fixed point along the flow. By combining numerical solutions of the flow equations using exact diagonalization and matrix-product operator-based techniques, as well as analytical approximations that are non-perturbative in the inverse drive frequency ($\Omega$), we have shown that freezing is approximate at any finite $\Omega$. Moreover, the violation of the emergent conservation law, quantified by the commutator $H_{0}(\lambda),\mathcal{Q}]$, vanishes as $1/\Omega^{2}$ in the high-frequency limit.
@@
One of the key contributions of this work is the analysis of thermalization dynamics of the system beyond the prethermal fixed point, where we find that the flow between the intermediate fixed points can be described using instanton events. For the sake of concreteness, we have chosen to study the driven spin chain as our main numerical example using complementary methods. However, the demonstrated features for freezing and thermalization are quite generic, as demonstrated by the beautiful agreement between our numerical results and the general model-independent analytical computations. Using a combination of different numerical diagnostics, we have demonstrated that the system flows to a random-matrix-like Hamiltonian when the frequency $\Omega$ is below a system-size-dependent threshold, $\Omega_{*}(L)$. Due to present numerical limitations of simulating only small system sizes, we have only been able to obtain the weakly increasing $L$-dependence of $\Omega_{*}(L)$ in a narrow range. Further numerical and analytical advances will be required to obtain the form of $\Omega_{*}(L)$ for a generic non-integrable Hamiltonian. If $\Omega_{*}(L)$ diverges with $L$ in the thermodynamic limit, that indicates fully thermalizing behavior even at dynamical freezing. However, it is possible that dynamical freezing may slow down the rate of divergence, or keep $\Omega_{*}(L)$ to be finite in the $L\rightarrow\infty$. This remains an interesting open problem for the future.
[…]
@@
We conclude by discussing several additional future directions. In the current work, we have only considered the cosine wave drive for simplicity, as there are just two Fourier components in the flow equation. Recently, there has been numerical evidence suggesting that dynamical freezing is more robust for square-wave drive [38, 36]. The flow formalism can be generalized to describe the square-wave drive, by introducing an infinite series of Hamiltonian components corresponding to the Fourier decomposition of the drive function [18]. Whether the flow away from the first prethermal fixed point is similar in character to the example studied here, but with possibly delayed timescales, is an interesting open question.
@@ Acknowledgements.
DC thanks Vedika Khemani for useful discussions, and Dan Mao for related collaborations. We thank Zihao Qi for discussions regarding the isospectral nature of the many-body spectrum during the flow. RM, HG and DC are supported in part by a New Frontier Grant awarded by the College of Arts and Sciences at Cornell University, and by a Sloan research fellowship from the Alfred P. Sloan foundation to DC. Some of the computations presented in this work are also supported in part by a NSF CAREER grant (DMR-2237522) to DC. HG is also supported by a Bethe-Wilkins-KIC postdoctoral fellowship at Cornell University.
DATA AVAILABILITY
All the theoretical data generated in this study have been deposited at Zenodo and are publicly available as of the date of publication: https://zenodo.org/records/18421731.
@@ Appendix A Connection between flow-renormalization and Schrieffer-Wolff transformation
Within the flow equation approach [55, 60, 91, 41, 15], the Hamiltonian $H$ is transformed continuously by defining a family of Hamiltonians $H(l)$, where $H(0)=H$ and $H(l)$ flows toward a diagonal (or block-diagonal) form as $l\to\infty$. The flow is governed by the differential equation,
@@
The ideas discussed above can be generalized to the Floquet context using the Sambe representation [67, 18]. In this representation, the time-dependent Hamiltonian ${H(t)}=\sum_{m}e^{im\Omega t}H_{m}$ is represented in an extended Hilbert space, where the Floquet operator $H(t)-i\partial_{t}$ is represented as
@@
Following Ref. [18], the Floquet flow is constructed as a unitary flow in the Sambe representation, which automatically preserves the eigenspectrum of $\mathcal{H}$} and hence the spectrum of $H_{F}$. To ensure unitarity, the flow should be generated by a commutator between the generator $\eta(l)$ and ${\mathcal{H}}(l)$ as in Eq. (29). Additionally, the generator should be chosen such that it strives to remove the oscillatory parts of $\mathcal{H}$} and it does not generate higher harmonics not present in $\mathcal{H}$}. The generator that satisfies the above requirements is
[…]
@@
Appendix B Additional numerical details
Here we provide some additional details about our numerical simulations. In our simulations, we did not explicitly fix a maximum bond dimension; instead, we used the default singular value cutoff in ITensors.jl, which is $10^{-12}$. For all numerical results involving MPO evolution, the bond dimensions of both $H_{0}(\lambda)$ and $H_{1}(\lambda)$ remained below 500. To compute the entanglement entropy, we initialized the MPS in a fully polarized product state along the $x$-direction and evolved it using the TDVP method. Since this initial state is unentangled, we employed the global-Krylov method to enrich the MPS with Krylov vectors. For the time evolution of the MPS, we employ the default single-site TDVP algorithm; for further details, see [53]. During the time evolution of the MPS, the SVD cutoff was set to $10^{-10}$. For the time evolution using the RK4 method, we set the integration step size to $10^{-2}$ for both the exact and MPS-based calculations. When studying the $\Omega$-scaling of the commutator at higher frequencies—where the associated time period becomes shorter—we reduce the step size to $10^{-3}$ to accurately resolve the faster dynamics.
Appendix C Additional numerical results
  • Effect of static transverse magnetic field:
    In continuation of the discussion in Sec. IV.2, in Fig. 10(a), we plot the expectation value of $S_{x}$ in the eigenbasis of $H_{0}(\lambda_{c})$, both with and without $B_{x}$, at and away from the freezing point. At the freezing point, the effective Floquet Hamiltonian to leading order is proportional to $YY+ZZ$ terms; see Eq. (70) in Appendix F. This structure gives rise to an emergent $\mathbb{Z}_{2}$ symmetry, resulting in degenerate Floquet eigenstates (equivalently, eigenstates of $H_{0}(\lambda_{c})$), and consequently, the expectation values of $S_{x}$ are nearly zero for all states. Introducing a small nonzero $B_{x}$ lifts the degeneracy, making the Floquet eigenstates to align with the eigenstates of $S_{x}$. In Fig. 10(b), we show the real-time magnetization profile for a thermal pure state evolving under $H_{0}(\lambda_{c})$.
    We find that the presence of $B_{x}$ enhances the quality of freezing. In Fig. 11 we show late (real-)time magnetization at the first freezing point for different frequency. Even though at the freezing point the dynamics “slows down” the magnetization still decays, indicating only an approximate conservation law. In this simulation, $\Omega$ is chosen large enough to avoid resonances between the eigenstates of the unperturbed part of the Hamiltonian. For smaller driving frequencies (but $\Omega\gtrsim J$), even at the first freezing point, the magnetization decays substantially. For the same driving frequency, at a higher-order freezing point (blue curve, $A=15.32$) compared with the lower-order case (yellow curve, $A=6.01$), the magnetization still decays from its initial value. This indicates that while the presence of a magnetic field enhances freezing at the level of the expectation value of $S_{x}$, it cannot prevent the eventual decay; see also the section on diagonal ensemble average in the Supplementary information for a related discussion. Whether this decay ultimately leads to thermalization to infinite temperature in the $L\rightarrow\infty$ limit remains an open question.
    Figure 10: (a) The values of $S_{x}$ for the eigenstates of $H_{0}(\lambda_{c})$ are plotted in decreasing order, comparing the freezing point (FP1) and the nonfreezing point (NFP), both with and without the magnetic field term in the $x$ direction. Without the magnetic field in the $x$ direction, at the freezing point, the leading-order effective Floquet Hamiltonian contains terms proportional to $YY+ZZ$, which implies a global $Z_{2}$ symmetry under $X_{i}\to-X_{i}$. As a result, the expectation value of $S_{x}$ in the eigenstates of $H_{0}(\lambda_{c})$ is vanishingly small. Parameters: $L=10$, $J=1$, $J_{2}=0.2$, $\Omega=5$, $\lambda_{c}=5.0$. (b) Magnetization is plotted as function of real time at the first freezing point with/without the magnetic field term $B_{x}$. Parameters: $L=10$, $J=1$, $J_{2}=0.5$, $\lambda_{c}=5.0$.
    Figure 11: Magnetization for a thermal pure state (evolving under Hamiltonian $H_{0}(\lambda_{c})$) as a function of real time at the first freezing point for two different driving frequencies. Parameters: $L=10$, $J=1$, $J_{2}=0.5$, $B_{x}=0.1$, $\lambda_{c}=10$.
    Figure 12: $P(\lambda\to\lambda_{c})$ for (L=10,20,40L=10,~20,~40) at the first freezing point (FP1) exhibits a universal $1/\Omega^{2}$ scaling. Inset: $Q(\lambda=\lambda_{c})$ for different system sizes exhibits an exponential decay with increasing driving frequency. The data near FP2 also follows the same frequency scaling.
  • Frequency scaling of $P(\lambda_{c})$:
    Recall the definition of $P(\lambda)$ in Eq. (12), which helps identify the extent of freezing and the emergent conservation law. In Fig. 12, we plot $P(\lambda_{c})$ at two distinct freezing points with increasing $\Omega$, where $\lambda_{c}$ is associated with the initial prethermal plateau. Interestingly, we find that $P(\lambda_{c})$ decays as $1/\Omega^{2}$ at large $\Omega$, which is consistent with our analytical results in Sec. VI. Relatedly, $Q(\lambda)$ (Eq. (13)) falls off exponentially with the drive frequency, $\Omega$ (Fig. 12 inset). We also point out that $P(\lambda)$, which is normalized to be an intensive quantity, already appears to show convergence to the thermodynamic limit for $L=20$, given the lack of any appreciable difference from the $L=40$ results.
  • Dynamical freezing over longer fRG times:
    In the main text, we showed results for the plateau associated with $P(\lambda)$ upto $\lambda\sim 1$. However, the plateau extends until later fRG times, as shown in Fig. 13, both at and away from freezing point for $\Omega=10$.
  • Renormalization of freezing points:
    To examine the possible renormalization of the freezing point away from the leading-order Floquet-Magnus predictions, we compute numerically the location of the minima of $P(\lambda_{c})$ for various driving frequencies. In Fig. 14, a fine-grid scan of $P(\lambda_{c})$ is performed near FP1 as a function of $A/\Omega$ for two different system sizes (L=8,10L=8,~10) using flow ED. We observe that with increasing frequency, the minimum of $P(\lambda_{c})$ approaches the value (=0.601=0.601) predicted by the zeroth-order Magnus expansion. However, for lower frequencies, there is a slight shift in the location of the freezing point.
    Figure 14: A fine-grid scan for $P(\lambda\to\lambda_{c})$ around FP1, for (L=8,10L=8,~10) and two different driving frequencies. The dashed line represents the prediction from the zeroth-order Magnus term.
  • Higher order freezing points:
    In Fig. 15, we illustrate the scenario where the driving frequency is comparable to the single-particle energy scales (in contrast to the choice of $\Omega$ in Fig. 3a), and the driving amplitude represents the largest energy scale in the system—effectively exploring higher-order freezing points while keeping the frequency fixed. In this case, we did not observe any qualitative differences; as the amplitude increases, the plateau value of $P(\lambda)$ decreases.
    Figure 15: $P(\lambda)$ is plotted as a function of RG time for different freezing points, where the value of the asymptotic value decreases continuously for higher order freezing points. Parameters: LL=10, JJ=1, $J_{2}=0.2$, $\Omega=2.0$.
  • Multiple separated instanton events:
    In Fig. 16(a), we plot the norm of $H_{1}$ as a function of fRG time using flow ED for a system of size $L=6$. The two well-separated instanton events are clearly visible, where the norms of $H_{1}(\lambda)$ peak at approximately $\lambda\approx 125$ and $\lambda\approx 380$, respectively. In Fig. 16(b), the corresponding norm of $H_{0}(\lambda)$ is shown, displaying a clear step-like feature precisely at the locations where $H_{1}(\lambda)$ peaks. Calculations for longer fRG times are constrained by our numerical precision, which in this case is double precision (binary64).
[…]
Appendix D Early-time analysis
@@
D.1 Review of Floquet-Magnus expansion
@@
$\begin{split}H_{\text{mov}}(t)&=W(t)^{\dagger}\left(H(t)-i\partial_{t}\right)W(t)\\ &=\exp\left(i\frac{2\sin(\Omega t){\mathrm{Ad}_{H_{1}}}}{\Omega}\right)H_{0}\,.\end{split}$ (36)
@@
and $I_{00}^{(1)}=-(i\pi)/(2\Omega)$. Since $h_{m},h_{-m}]=0$, we only need $I_{m0}^{(1)}$ and $I_{0m}^{(1)}$. Therefore we obtain
@@
D.2 Solution of flow equation
@@
Here $\bar{H}_{1}(\lambda)$ and $\bar{H}_{1}(\lambda)^{\dagger}$ depend on $H_{0}(\lambda)$, as described by Eqs. (47a), and (47b), and $\{\cdot,\cdot\}$ denotes the usual anticommutator.
We emphasize that all of the manipulations thus far are exact. To make connection with the Floquet-Magnus expansion above, we can expand Eq. (D.2) in powers of $H_{0}(0)$, i.e.
@@
  • First order term: We first solve for $H_{0}^{(1)}$, which is obtained by setting $\bar{H}_{1}(\lambda)=\bar{H}_{1}(\lambda)^{\dagger}=H_{1}(0)$ in Eq. (D.2), and the associated linear second-order differential equation:
@@
    - $J_{0}\bigg(\frac{A}{\Omega}\bigg)Y_{1}\bigg(\frac{A}{\Omega}e^{-\Omega\lambda}\bigg)\bigg].$
@@
  • Second order term: The term, $H_{0}^{(2)}(\lambda)$, can be obtained by expanding Eq. (D.2) to ${\cal{O}}(\zeta^{2})$, with a resulting equation:
@@
    The source term, $S^{(2)}(\lambda)$, is obtained by expanding the last term of Eq. (D.2) to ${\cal{O}}(\zeta^{2})$,
@@
    $+\bigg\{{\mathrm{Ad}}_{\bar{H}_{1}^{(0)}(\lambda)},{\mathrm{Ad}}_{\bar{H}_{1}^{(1)}(\lambda)^{\dagger}}\bigg\}\bigg]H_{0}^{(1)}(\lambda)\,.$ (55)
@@
    Interestingly, due to conjugation, $\bar{H}_{1}^{(1)}(\lambda)^{\dagger}=-\bar{H}_{1}^{(1)}(\lambda)$, the two terms in Eq. (D.2) cancel, and hence the source term, $S^{(2)}(\lambda)=0$, in Eq. D.2 vanishes. Combining with the initial condition $H_{0}^{(2)}(0)=\partial_{\lambda}H_{0}^{(2)}(0)=0$, the second order term vanishes identically.
@@
Appendix E Estimate of thermalization timescales